# One Demo Is All It Takes: Planning Domain Derivation with LLMs from A Single Demonstration

### v1

@@ 6 Experiments and Baselines
Our Experiments are conducted in PyBullet Simulation (Coumans and Bai, 2021), with the symbolic solver from PDDLStream (Garrett et al., 2020) and the motion planning algorithm from the PyBullet-Planning (Garrett et al., 2015; Garrett, 2018). The LLM we used is GPT-4o (OpenAI, 2023).
Task Diversity: To ensure robustness and broad applicability, we sampled over 1,200 planning tasks across nine distinct environments, including stacking, unstacking, rearrangement, alignment, parts assembly, color classification, burger cooking, and bridge building. The details of the tasks are provided in Section A.4. Each task category includes tasks of 3 to 20 objects, with 10 distinct tasks sampled for each object count. The resulting task plan lengths ranged from 6 to 510 steps, reflecting a wide spectrum of planning horizons. Our experiments spanned over 150 unique predicates.
Task Complexity: In task design, we consider two types of complexity: domain derivation complexity and planning complexity. Domain derivation complexity is determined by the number of predicates imagined and actions invented by PDDLLM; the more predicates and actions, the higher the complexity. Planning complexity is influenced by both the planning domain and the task. Given $n$ objects, a task plan of length $l$, and $m$ actions in the domain, the branching factor at each step is $m\times n$, resulting in a approximate complexity of $(m\times n)^{l}$.
Table 1 showcase the planning complexity (in approximate order of magnitude) of the most difficult problem in the category and the domain derivation complexity. The Tower of Hanoi task exhibits the greatest planning complexity, while bridge building presents the highest domain derivation complexity. In contrast, stack and unstack are simpler in both complexity measures.
Task Stack Unstack Color Classify Alignment Parts Assembly Rearrange Burger Cook Bridge Build Tower of Hanoi
Max Planning Complexity $10^{58}$ $10^{58}$ $10^{71}$ $10^{58}$ $10^{58}$ $10^{137}$ $10^{48}$ $10^{36}$ $10^{307}$
Domain Derivation Complexity 90 90 111 94 92 96 114 128 94
@@ 6.1 Baselines and Ablations
We implement six baselines and one ablation of PDDLLM to comprehensively evaluate our method. GPT-4o is used as the default LLM unless otherwise specified, and the motion planning algorithm is kept the same for all methods. The complete details are provided in Section A.2
@@
  • LLMTAMP-FF: Following the method by Huang et al. (2023); Chen et al. (2023), LLMTAMP-FF extends LLMTAMP with a failure feedback loop.
@@
Robot planning is required to be real-time in robot deployment, imposing constraints on the planning time allowance (Gammell et al., 2015; Garrett et al., 2015, 2020). In our experiment, a uniform planning time limit of 50 seconds is applied to all planning problems and methods. We measure performance using the planning success rate, as in other studies (Huang et al., 2024; Silver et al., 2023; Kumar et al., 2023). The planning time and token cost are used to measure the planning cost (Zhong et al., 2024). Three parallel runs were conducted to compute the mean and standard error of the means for the planning success rate.
@@
Method Expert LLMTAMP LLMTAMP-FF LLMTAMP-FR RuleAsMem PDDLLM
Stack $96.1\pm 0.2$ $41.7\pm 4.3$ $70.8\pm 1.4$ $64.2\pm 3.1$ $85.5\pm 2.9$ 97.5 $\pm$ 1.6
Unstack $96.1\pm 0.2$ $89.4\pm 1.5$ $94.6\pm 0.9$ $92.1\pm 2.3$ $88.4\pm 1.2$ 97.7 $\pm$ 0.7
@@ 7 Results
Through the experiments, we aim to answer the following research questions: (1) How does PDDLLM perform relative to other LLM-based planners? (2) Can PDDLLM generalize to unseen, more complex tasks? (3) Does PDDLLM derive high-quality domains with performance comparable to expert designs? (4) How does PDDLLM’s token cost compare to other LLM-based planners?
@@
Table 2 presents the planning success rate of all evaluated methods across all tasks, measured with a 50-second time limit. PDDLLM shows clear advantages in planning efficiency and generalizability over baseline methods. While LLM-based baselines perform competitively in simpler tasks like stacking and unstacking, their performance drops sharply in complex tasks such as rearrangement, burger cooking, bridge building, and Tower of Hanoi. In contrast, PDDLLM maintains strong performance across all task categories, achieving an over 40% improvement in overall planning success rate compared to the best LLM-based planner baseline (i.e., LLMTAMP-FF). Even the ablated variant of PDDLLM, RuleAsMem, outperforms the LLM-based planners by at least 17.4% in overall success rate. Compared to PDDLLM, RuleAsMem exhibits less stability. It performs well in simpler tasks but struggles in more complex ones, suggesting the LLMs struggle to understand complex domains. In addition to the results evaluated with a 50-second time window, we also report the overall success rates of the main methods across varying time limits. As shown in Figure 3(right), PDDLLM consistently outperforms the baseline methods and demonstrates superior planning performance across all time limits. Notably, it reaches performance saturation the fastest, highlighting its superior time efficiency among all evaluated approaches.
We further compare PDDLLM’s planning ability with more powerful reasoning LLMs (OpenAI-o1 and DeepSeek-R1) in Table 4. While reasoning-based models exhibit strong planning capabilities, their high computational cost prevents them from completing within the 50-second window. To sufficiently compare the planning capabilities, we extend the time limit to 500 seconds and evaluate them on the most challenging tasks. Although o1-TAMP and R1-TAMP show remarkable improvement compared to GPT-4o-based LLMTAMP, they remain less robust in long-horizon planning and fail to match PDDLLM’s performance. In contrast, our method, relying solely on GPT-4o, consistently achieves higher success rates across complex tasks. Detailed time costs for reasoning models are provided in Section A.6.
@@
The experimental results highlight PDDLLM’s ability to handle increasing complexity in both planning and domain derivation, consistently outperforming baseline methods. In terms of planning complexity, as shown in Figure 3(left), PDDLLM maintains robust planning performance as task complexity grows, achieving high success rates even in scenarios involving up to 20 objects. As shown in Table 2, performance degradation is observed in more challenging tasks such as rearrangement, where longer action sequences are required and there are more complex motion constraints. From the perspective of domain derivation complexity, PDDLLM remains effective even in tasks demanding the generation of over 100 predicates. However, success rates drop in the most complex domains, such as bridge building, primarily due to missing supporting predicates. More details of the generalization of each task is provided in section A.5.
PDDLLM demonstrates a strong ability to integrate knowledge across demonstrations. The modular nature of PDDL action syntax allow the easy transfer of actions learned from different demonstrations among domains. Despite receiving only one example each for stacking and unstacking, it successfully combines the “stack” action and the “unstack” action to solve rearrangement tasks. Similarly, in the bridge building domain, PDDLLM successfully combines the “align” action and the “stack” action using one demonstration of cube alignment and one of stacking. The high success rate of these tasks in Table 2 underscore the robustness of the derived planning domains.
@@
We evaluate the quality of planning domains generated by PDDLLM by comparing the percentage of missing or redundant predicates and actions, as well as the planning success rate across a range of tasks, against expert-designed domains. Table 3 reports the percentage of missing or redundant elements in the final domains; tasks with no such issues are omitted. These results indicate that PDDLLM produces high-quality domains with minimal errors, even in complex scenarios. While a few predicates may be absent, the overall logical structure remains sound, as reflected in the consistently high planning success. As shown in Table 2, the derived domains achieve a success rate of 93.3%, closely matching the performance of expert-crafted domains.
Table 3: Percentage of missing or redundant predicates and actions across selected domains.
Task Stack Burger Cooking Bridge Building Tower of Hanoi
Percentage of Missing predicates or actions 0.0% 18.8% 16.7% 0.0%
Percentage of redundant predicates or actions 10.0% 0.0% 0.0% 11.1%
@@
We compare the token efficiency of our method against LLM-based baselines on the three most complex tasks, as shown in Table 4. Although the GPT-4o-based LLMTAMP incurs lower total token costs, it performs poorly across all three tasks. Compared to o1-TAMP and R1-TAMP, our PDDLLM uses significantly fewer tokens while consistently achieving better performance. These results also underscore the challenge of deploying reasoning LLMs on real robot systems, given their high token consumption. Beyond the main research questions, we also evaluated the robustness of our method under variations in prompting styles. Results show that PDDLLM remains stable across different prompts. Full details are provided in Section A.7.
@@ 8 Real Robot Deployment
Figure 4: Real-robot experiment: A Piper arm making two burgers
We evaluate PDDLLM’s validity on real-world robotic systems to demonstrate its cross-platform deployability. The system is tested on both the Agilex Piper Arm and the Franka Panda Arm. ArUco 5×5 markers are used for pose estimation, while ROS2 and MoveIt2 serve as the robot driver and motion planning framework. Both robots successfully complete real-world planning tasks through the direct deployment of PDDLLM, demonstrating its effectiveness across different hardware platforms and validating its applicability in real-world robotic scenarios. Four real robot experiments, including stacking, bridge building, cooking burgers, and the tower of Hannoi are performed. Full information can be found in Section A.1. Figure 4 showcases the Piper arm making burgers, with the task and motion plan from PDDLLM.
@@ 9 Limitation
A current limitation of PDDLLM lies in its occasional omission of logically complex predicates, which contributes to the observed decline in success rates for more complex domains such as Burger Cooking and Bridge Building, as shown in Table 2. For example, in the bridge building task, the robot must assemble the top surface only after all base components are correctly placed. If the surface is assembled prematurely, it becomes difficult to generate feasible manipulation trajectories for the remaining base components, often resulting in planning failure. In contrast, the expert-designed planning domain addresses this issue by introducing a predicate (all_base_finished) as a precondition for surface assembly, thereby enforcing the correct ordering of actions. Recent research has begun to address this limitation, for example, by leveraging VLMs to invent new predicates (Liang et al., 2024), or by refining the domain through environmental feedback (Huang et al., 2024).
@@ 10 Conclusion
This paper presents PDDLLM, the first approach in the field to generate a complete planning domain from scratch, without relying on any predefined predicates or actions. By extracting logical structures directly from pre-trained LLMs, PDDLLM autonomously derives both predicates and actions, enabling fully automated domain construction. Evaluated across a wide range of environments, PDDLLM demonstrates high quality in domain derivation and strong generalizability across diverse task categories. Moreover, when integrated with the LoCAS framework, PDDLLM fully automates the integration between the PDDL planning domain and the low-level motion planner. This level of automation significantly improves usability and positions the framework as an adaptable and scalable solution for robotic planning and decision-making. Compared to existing methods, PDDLLM outperforms other LLM-based baselines and closely matches the performance of expert-designed planning domains, particularly in complex and long-horizon planning scenarios. Future work will focus on integrating perception into the system, enabling domain derivation from raw sensory inputs. From the aspect of broader impact, PDDLLM lowers the barrier to deploying robots by enabling domain derivation from demonstrations, making automation more accessible and adaptable.
@@ A.6 Time Cost of LLMTAMP Reasoning Model Variants
As shown in Table 5, our method consistently yields lower average planning times across all tasks compared to the LLMTAMP reasoning model variants. This improvement is primarily attributed to PDDLLM’s ability to structurally summarize the reasoning process into a standardized planning domain during a one-time offline inference step. As a result, no additional reasoning is required at test time. In contrast, LLMTAMP variants conduct reasoning independently for each task instance, leading to significantly higher computational costs. Such overhead makes these models impractical for deployment on physical robots, where real-time planning capabilities are often essential. As shown in Table 4, while o1-TAMP achieves comparable performance in certain tasks, the substantially higher time cost undermines its practical value, particularly in real robotic scenarios.
@@ A.7.1 Parallel prompting for predicate imagination
In this prompt variation, we perform parallel prompting of the LLM multiple times to obtain diverse responses simultaneously. These outputs are subsequently analyzed and aggregated to synthesize an optimal solution. While the core prompting procedure follows the same structure as described in Section A.3.1, additional post-processing steps are applied to summarize and consolidate the results. For prompt naming, we further prompt the LLM to select the most suitable name from among the parallel outputs. The prompt template is shown as the following:
@@ A.7.2 Altering the prompting sequence of simulation outcomes
In this prompt variation, we alter the ordering of dimensions presented to the LLM. While the prompt template remains similar to that described in Section A.3.1, we do not adhere to a fixed dimension sequence. Instead, we shuffle the order of dimensions within the prompt to evaluate whether this affects the generated output. Some examples are provided here.
@@ A.7.3 Template tuning
The results, presented in Table 6, demonstrate that our method remains robust across different prompting styles. Regardless of the prompt variation, the generated planning domains consistently solve the test tasks with a success rate approaching 100%. Minor fluctuations are attributed to randomness in the planning search process and occasional motion execution failures.

### v3

@@ 6 Experiments and Baselines
Building on evaluation practices in TAMP, we consider tasks with concrete conditions such as object position, orientation, size, and color (Silver et al., 2023; Garrett et al., 2020; Kumar et al., 2023; Liang et al., 2024; Huang et al., 2025). We structure our experiments along three dimensions, Task Diversity, Task Complexity, and Knowledge Transferability, whose definitions are provided in Section˜A.10.1, with task details described in Section˜A.4. Our experiments are conducted in PyBullet Simulation (Coumans and Bai, 2021), with the symbolic solver from PDDLStream (Garrett et al., 2020) and the motion planning algorithm from PyBullet-Planning (Garrett et al., 2015; Garrett, 2018). The LLM we used is GPT-4o (OpenAI, 2023).
@@ 6.1 Baselines and Ablations
We implement six baselines and one ablation of PDDLLM to comprehensively evaluate our method. GPT-4o is used as the default LLM unless otherwise specified, and the motion planning algorithm is kept the same for all methods. The complete details are provided in Section˜A.2
@@
  • LLMTAMP-$F$: Following the method by Huang et al. (2023); Chen et al. (2023), LLMTAMP-$F$ extends LLMTAMP with a failure feedback loop.
@@
Robot planning is required to be real-time in robot deployment, imposing constraints on the planning time allowance (Gammell et al., 2015; Garrett et al., 2015, 2020). In our experiment, a uniform planning time limit of 50 seconds is applied to all planning problems and methods. We measure performance using the planning success rate, as in other studies (Huang et al., 2025; Silver et al., 2023; Kumar et al., 2023). The planning time and token cost are used to measure the planning cost (Zhong et al., 2024). Three parallel runs were conducted to compute the mean and standard error for the planning success rate.
@@
Method Expert LLMTAMP LLMTAMP-$F$ LLMTAMP-FR RuleAsMem PDDLLM
Stack $98.5\pm 0.8$ $41.7\pm 4.3$ $70.8\pm 1.4$ $64.2\pm 3.1$ $85.5\pm 2.9$ 97.5 $\pm$ 1.6
Unstack $100\pm 0.0$ $89.4\pm 1.5$ $94.6\pm 0.9$ $92.1\pm 2.3$ $88.4\pm 1.2$ 97.7 $\pm$ 0.7
@@ 7 Results
Through the experiments, we aim to answer the following research questions: (1) How does PDDLLM perform relative to other LLM-based planners? (2) Can PDDLLM generalize to unseen, more complex tasks? (3) Does PDDLLM derive high-quality domains with performance comparable to expert designs? (4) How does PDDLLM’s token cost compare to other LLM-based planners? (5) How stable is PDDLLM when faced with suboptimal demonstration?
@@
Table˜1 presents the planning success rate of all evaluated methods across all tasks, measured with a 50-second time limit. PDDLLM shows clear advantages in planning efficiency and generalizability over baseline methods. While LLM-based baselines perform competitively in simpler tasks like stacking and unstacking, their performance drops sharply in complex tasks such as rearrangement, burger cooking, bridge building, and Tower of Hanoi. In contrast, PDDLLM maintains strong performance across all task categories, achieving an over 40% improvement in overall planning success rate compared to the best LLM-based planner baseline (i.e., LLMTAMP-FF). Even the ablated variant of PDDLLM, RuleAsMem, outperforms the LLM-based planners by at least 17.4% in overall success rate. Compared to PDDLLM, RuleAsMem exhibits less stability. It performs well in simpler tasks but struggles in more complex ones, suggesting the LLMs struggle to understand complex domains. In addition to the results evaluated with a 50-second time window, we also report the overall success rates of the main methods across varying time limits. As shown in Figure˜3(right), PDDLLM consistently outperforms the baseline methods and demonstrates superior planning performance across all time limits. Notably, it reaches performance saturation the fastest, highlighting its superior time efficiency among all evaluated approaches.
We further compare PDDLLM’s planning ability with more powerful reasoning LLMs (OpenAI-o1 and DeepSeek-R1) in Table˜2. While reasoning-based models exhibit strong planning capabilities, their high computational cost prevents them from completing within the 50-second window. To sufficiently compare the planning capabilities, we extend the time limit to 500 seconds and evaluate them on the most challenging tasks. Although o1-TAMP and R1-TAMP show remarkable improvement compared to GPT-4o-based LLMTAMP, they remain less robust in long-horizon planning and fail to match PDDLLM’s performance. In contrast, our method, relying solely on GPT-4o, consistently achieves higher success rates across complex tasks. Detailed time costs for reasoning models are provided in Section˜A.6.
@@
The experimental results highlight PDDLLM’s ability to handle increasing complexity in both planning and domain derivation, consistently outperforming baseline methods. In terms of planning complexity, as shown in Figure˜3(left), PDDLLM maintains robust planning performance as task complexity grows, achieving high success rates even in scenarios involving up to 20 objects. As shown in Table˜1, performance degradation is observed in more challenging tasks such as rearrangement, where longer action sequences are required and there are more complex motion constraints. From the perspective of domain derivation complexity, PDDLLM remains effective even in tasks demanding the generation of over 100 predicates. However, success rates drop in the most complex domains, such as bridge building, primarily due to missing supporting predicates. More details of the generalization of each task are provided in section˜A.5.
PDDLLM demonstrates a strong ability to integrate knowledge across demonstrations. The modular nature of PDDL action syntax allows the easy transfer of actions learned from different demonstrations among domains. Despite receiving only one example each for stacking and unstacking, it successfully combines the “stack” action and the “unstack” action to solve rearrangement tasks. Similarly, in the bridge building domain, PDDLLM successfully combines the “align” action and the “stack” action using one demonstration of cube alignment and one of stacking. The high success rate of these tasks in Table˜1 underscores the robustness of the derived planning domains.
@@
We evaluate the quality of planning domains generated by PDDLLM by comparing the percentage of missing or redundant predicates(actions) and the planning success rate against expert-designed domains. The full definition of missing or redundant can be found in Section˜A.11. None of the generated domains are missing actions. The predicate statistics are reported in Table˜4. The table summarizes the percentage of missing or redundant predicates, omitting tasks without such issues. These results show that PDDLLM produces high-quality domains with minimal errors, even in complex scenarios. While a few predicates may be absent, the overall logical structure remains sound, as reflected in the high overall planning success rate of 93.3% shown in Table˜1, closely matching the performance of expert-crafted domains.
@@
We compare the token efficiency of our method against LLM-based baselines on the three most complex tasks, as shown in Table˜2. Although the GPT-4o-based LLMTAMP incurs lower total token costs, it performs poorly across all three tasks. Compared to o1-TAMP and R1-TAMP, PDDLLM uses significantly fewer tokens while consistently achieving better performance. These results also underscore the challenge of deploying reasoning LLMs on real robot systems, given their high token consumption. Beyond the main research questions, we also evaluated the robustness of our method under variations in prompting styles. Results show that PDDLLM remains stable across different prompts. Full details are provided in Section˜A.7.
@@
Q5. Suboptimal demonstrations:
Lastly, we conducted experiments on the bridge-building task to evaluate how PDDLLM adapts to sub-optimal demonstrations by intentionally masking or perturbing input trajectories. The full skill set required for this task consists of pick, stack, and align actions. In the first experiment, only a pick trajectory was provided; subsequent demonstrations incrementally added stack and then align. As shown in Table˜4, success rates improved as demonstrations covered more required actions, reflecting the system’s ability to aggregate knowledge across trajectories. In a fourth experiment, we introduced a noisy trajectory corresponding to an unrelated unstack action. Although this action was initially incorporated into the domain, its negative impact was light: it was never invoked in a solution and was later pruned due to low use frequency.
Table 3: Percentage of missing or redundant predicates and actions across tasks.
Task Missing Redundant
Stack 0.0% 12.5%
Burger Cooking 22.2% 0.0%
Bridge Building 22.2% 0.0%
Tower of Hanoi 0.0% 14.3%
Table 4: Bridge-building success rate (%) under varying demonstration conditions.
Class Demonstration Type Success Rate (%) $\uparrow$
Missing pick $0.0\pm 0.0$
Missing pick & stack $20.0\pm 0.0$
Complete pick & stack & align $86.7\pm 3.8$
Redundant pick & stack & align & unstack $83.3\pm 3.3$
@@ 8 Real Robot Deployment
Figure 4: Real-robot experiment in three different platforms
We validate PDDLLM on real robots to demonstrate cross-platform deployability. The system is tested on the Agilex Piper, Franka Panda, and the UR5e, using ArUco markers for pose estimation and ROS2 for control. All platforms successfully complete tasks, including stacking, bridge building, burger cooking, and the Tower of Hanoi through direct deployment of PDDLLM (details in Section˜A.1). We also tested how PDDLLM integrates with learned skills such as VLA: logical actions serve directly as prompts to condition policies, bypassing explicit motion constraints. A real-robot experiment with VLA demonstrates this seamless interface (details in Section˜A.1.3), highlighting the framework’s flexibility across classical and learning-based control.
@@ 9 Limitation
A limitation of PDDLLM is its occasional omission of complex predicates, leading to lower success rates in harder domains like Burger Cooking and Bridge Building (Table˜1). For example, in bridge building the surface may be assembled before the base is complete, making subsequent placements infeasible. Expert-designed domains address this by introducing a predicate (all_base_finished) to enforce ordering. Recent work has sought to overcome this limitation by performing domain refining (Liang et al., 2024; Huang et al., 2025). As with many TAMP methods, modeling complex geometric conditions such as deformable objects is challenging in logical space. Nevertheless, we demonstrate that PDDLLM can be integrated with approaches targeting such problems, including VLA methods.
@@ 10 Conclusion
This paper presents PDDLLM, the first approach in the field to generate a complete planning domain from scratch, without relying on any predefined predicates or actions. By extracting logical structures directly from pre-trained LLMs, PDDLLM autonomously derives both predicates and actions, enabling fully automated domain construction. Evaluated across a wide range of environments, PDDLLM demonstrates high quality in domain derivation and strong generalizability across diverse task categories. Moreover, when integrated with the LoCA framework, PDDLLM fully automates the integration between the PDDL planning domain and the low-level motion planner. This level of automation significantly improves usability and positions the framework as an adaptable and scalable solution for robotic planning and decision-making. Compared to existing methods, PDDLLM outperforms other LLM-based baselines and closely matches the performance of expert-designed planning domains, particularly in complex and long-horizon planning scenarios. Future work will focus on integrating perception into the system, enabling domain derivation from raw sensory inputs. From the aspect of broader impact, PDDLLM lowers the barrier to deploying robots by enabling domain derivation from demonstrations, making automation more accessible and adaptable.
@@ A.1.2 Agilex Piper Arm
A.1.3 VLA Experiment with UR5e
Figure 9: UR5e Arm Stacking Tabletop Objects
Experiment with the vision-language-action model is conducted using a $\pi$0 model  (Black et al., 2024), finetuned on 200 real-world pick and place trajectories. The resulting policy serves as the action module within our pipeline and is invoked by the PDDLLM planning system. In this setup, PDDLLM generates a high-level task plan, while the execution of each logical action in the plan is carried out directly by the Pi0 policy. Importantly, the VLA experiment demonstrates the flexibility of our framework: the action execution layer can be implemented with either a classical motion planner or a learning-based policy, depending on the task requirements.
@@ A.6 Time Cost of LLMTAMP Reasoning Model Variants
As shown in Table˜5, our method consistently yields lower average planning times across all tasks compared to the LLMTAMP reasoning model variants. This improvement is primarily attributed to PDDLLM’s ability to structurally summarize the reasoning process into a standardized planning domain during a one-time offline inference step. As a result, no additional reasoning is required at test time. In contrast, LLMTAMP variants conduct reasoning independently for each task instance, leading to significantly higher computational costs. Such overhead makes these models impractical for deployment on physical robots, where real-time planning capabilities are often essential. As shown in Table˜2, while o1-TAMP achieves comparable performance in certain tasks, the substantially higher time cost undermines its practical value, particularly in real robotic scenarios.
@@ A.7.1 Parallel prompting for predicate imagination
In this prompt variation, we perform parallel prompting of the LLM multiple times to obtain diverse responses simultaneously. These outputs are subsequently analyzed and aggregated to synthesize an optimal solution. While the core prompting procedure follows the same structure as described in Section˜A.3.1, additional post-processing steps are applied to summarize and consolidate the results. For prompt naming, we further prompt the LLM to select the most suitable name from among the parallel outputs. The prompt template is shown as the following:
@@ A.7.2 Altering the prompting sequence of simulation outcomes
In this prompt variation, we alter the ordering of dimensions presented to the LLM. While the prompt template remains similar to that described in Section˜A.3.1, we do not adhere to a fixed dimension sequence. Instead, we shuffle the order of dimensions within the prompt to evaluate whether this affects the generated output. Some examples are provided here.
@@ A.7.3 Template tuning
The results, presented in Table˜6, demonstrate that our method remains robust across different prompting styles. Regardless of the prompt variation, the generated planning domains consistently solve the test tasks with a success rate approaching 100%. Minor fluctuations are attributed to randomness in the planning search process and occasional motion execution failures.